OpenStudy (anonymous):

Find the missing terms of ______, 4/9, ______, ______, ______, 22/9.

OpenStudy (anonymous):

I'm stuck. All I've figured out is 2/1 from 22/9 - 4/9.

OpenStudy (anonymous):

It is an arithmetic series. Would I use the formula somehow?

OpenStudy (campbell_st):

well there 4 terms and a distance of 18 so its 4.5 is added (4 - 4.5/9, 4/9,(4 + 4.5)/9 (4 + 4.5+4.5)/9 , (4 + 4.5 +4.5+4.5)/9, 22/9 just simplfy the fractions.

OpenStudy (anonymous):

\[-\frac{ 1 }{ 18 }\]\[\frac{ 4 }{ 9 }\]\[\frac{ 17 }{ 18 }\]\[\frac{ 35 }{18 }\]\[\frac{ 22 }{ 9 }\]

OpenStudy (anonymous):

They can't be decimals, so is that about an approximate?

OpenStudy (campbell_st):

no as fractions you have exact values... and changing 8.5/9 to 17/18 is the correct process...

OpenStudy (anonymous):

Hm, okay. Oh I missed one. There's also 13/9, I think.